# Aligning Individual and Collective Objectives in Multi-Agent Cooperation

### v2

@@ 1 Introduction
Recent studies in mixed-motive cooperative MARL have largely employed hand-crafted designs to promote collaboration, especially with focus on automatically modifying rewards. One popular approach is to align objectives as per existing mechanisms in cooperative games, such as reputation (Anastassacos et al., 2021), norms (Vinitsky et al., 2023), and contracts (Hughes et al., 2020). Another prevalent method leverages intrinsic motivation to align individual and collective objectives, enhancing altruistic collaboration by integrating heuristic knowledge into the incentive function. Conventionally, some works accumulate individual rewards with the group to promote altruistic conduct (Hostallero et al., 2020; Peysakhovich and Lerer, 2018; Apt and Schäfer, 2014; Roesch et al., 2024). Furthermore, some studies derive more sophisticated preference signals from the rewards of other agents (Hughes et al., 2018; McKee et al., 2020a). Additionally, several approaches aim to learn the potential influences of an agent’s actions on others (Yang et al., 2020; Jaques et al., 2019; Lu et al., 2022). Most of these algorithms rely heavily on carefully crafted designs, necessitating significant human expertise and detailed domain knowledge. On the other hand, several studies focus on automatically modifying rewards by learning additional weights to adjust the original objectives (Gemp et al., 2020; Kwon et al., 2023). However, these studies often suffer from a lack of interpretability and insufficient theoretical analysis regarding the alignment of individual and collective objectives.
@@ 2 Related Work
Mixed-motive cooperation refers to scenarios where the group’s objectives are sometimes aligned and at other times conflicted (Philip S. Gallo and McClintock, 1965). Recently, there has been a surge in academic interest in the Sequential Social Dilemma (SSD) (Leibo et al., 2017), which expands the concept from its roots in matrix games (Macy and Flache, 2002) to Markov games. Prosocial (Peysakhovich and Lerer, 2018) improves collective performance by blending individual rewards to redraft the agent’s overall utility. While PED-DQN enables agents to incrementally adjust their reward functions for enhanced collaborative action through inter-agent evaluative signal exchanges (Hostallero et al., 2020), Gifting directly reward other agents as part of action space (Lupu and Precup, 2020). Inequity aversion further integrates the concept into Markov games by adding envy (disadvantageous inequality) and guilt (advantageous inequality) rewards to the original individual rewards (Hughes et al., 2018). Simultaneously, Social Value Orientation (SVO) strategy introduces a unique shared rewards-based compensation approach, encouraging behavior modifications in line with interdependence theory (McKee et al., 2020b). The LIO strategy bypasses the need to modify extrinsic rewards by empowering an agent to directly influence its partner’s actions (Yang et al., 2020). Meanwhile, other studies introduce new cooperative mechanisms such as incorporating a reputation model (McKee et al., 2023), social norms (Vinitsky et al., 2023). Recently, Roesch et al. (2024) proposed a selfishness level method that incorporates social welfare into individual rewards to enhance altruistic cooperation. Additionally, several works have explored automatically modifying rewards online. D3C (Gemp et al., 2020) learns to mix rewards to improve efficiency in a Nash equilibrium. Kwon et al. (2023) extended D3C by addressing the problem of automatically modifying individual agent objectives to optimize a desired global objective. Despite these advances, many current studies lack both cost efficiency in design and theoretical analysis of alignment and convergence. To address this, our study applies a gradient perspective to align goals and further investigates the learning dynamics of our proposed method.
@@ 4.1 Differentiable Mixed-motive Game
Consider a two-player DMG with $\ell_{1}(a_{1},a_{2})=-sin(a_{1}a_{2}+a_{2}^{2})$ and $\ell_{2}(a_{1},a_{2})=-[cos(1+a_{1}-(1+a_{2})^{2})+a_{1}a_{2}^{2}]$. The rewards for the two players are the negation of their respective losses.
@@ 4.2 Altruistic Gradient Adjustment
Proposition 4.2 (Altruistic Gradient Adjustment).
@@
Note that the Hessian matrix ${\bm{H}}_{c}$ is symmetric. For brevity, we denote $\nabla{\mathcal{H}}_{c}$ as $\nabla_{\bm{w}}\left(\frac{1}{2}\|{\bm{\xi}}_{c}\|^{2}\right)$, which simplifies to ${\bm{H}}_{c}^{T}{\bm{\xi}}_{c}$. In our AgA method, the formula $\xi_{c}+\lambda(\xi+{\bm{H}}^{T}_{c}\xi_{c})$ includes the Hessian matrix, but we do not compute it directly. Instead, we calculate Hessian-vector products ${\bm{H}}^{T}_{c}\xi_{c}$, which suffice for determining the adjustment gradient and the sign of $\lambda$. This approach reduces computational complexity, enhancing the effectiveness of the AgA method. The cost for computing Hessian-vector products ${\bm{H}}^{T}_{c}\xi_{c}$ is $\mathcal{O}(n)$ for $n$ weights (Pearlmutter, 1994).
@@ 5 Experiments
To assess the effectiveness of the proposed AgA algorithm, we perform comprehensive experiments across various environments, ranging from simple to complex, and from small to large scales. The initial environment involves a two-player public goods game (see Section 5.1), followed by sequential social dilemma environments (see Section 5.2): Cleanup and Harvest, involving 5 homogeneous players. The final test is conducted in our specially developed selfish-MMM2 environment (see Section 5.3), which is both more complex and larger in scale, featuring approximately 10 controlled agents of 3 distinct types, competing against 11 adversaries. Comparative experiments employ commonly used baseline approaches, such as simultaneous optimization with individual and collective losses (termed Simul-Ind and Siml-Co), CGA (Mescheder et al., 2017), SGA (Balduzzi et al., 2018), SVO (McKee et al., 2020a), and the selfishness level approach (abbreviated as SL) (Roesch et al., 2024).
@@ 5.1 Two-Player Public Goods Game
Table 1: The comparison of the average individual rewards (denoted as $r_{1},r_{2}$), social welfare (denoted as $SW$) on two- player public goods game. We show the mean of rewards and 95% confidence interval utilizing 50 random runs.
@@
Results. Table 1 presents a comparison of individual rewards and the collective outcome, derived from 50 random trials, each limited to 100 steps. The rows $r_{1},r_{2}$ and $SW$ represent the rewards accrued by each individual player and the collective group, respectively. The social optimum of the game occurs when all players allocate their entire budget, achieving the highest possible social welfare score of $3$. Among all algorithms, AgA achieves the closest social welfare score to this optimum, with a value of 2.90. Furthermore, AgA exhibits the greatest degree of fairness, with two players receiving nearly identical rewards, in contrast to the baseline algorithms. Furthermore, Fig. 4 in Appendix D illustrates the action distributions of these algorithms, showing that AgA actions are more tightly clustered around the social optimum actions $(1,1)$. This further indicates that AgA approaches the social optimum more effectively.
@@
5.2 Sequential Social Dilemma: Cleanup and Harvest
Baseline Comparison. As illustrated in Fig. 3(g) and Fig. 3(h), AgA demonstrates a notable improvement in social welfare over Simul-Ind, Simul-Co, CGA, SVO, and SL. Particularly, within the Cleanup scenario, AgA recorded an average social welfare of 105.15, marking approximately a 56% increase over the runner-up method SL. In the Harvest scenario, AgA shows an average improvement of 33.55 in social welfare over the next best method SVO.
@@ 6 Conclusion
Limitations. Despite conducting a series of experiments to verify the proposed AgA algorithm, including using a new large-scale mixed-motive cooperation testbed, these environments remain somewhat removed from real-world applications. Moreover, while we do not explicitly derive the Hessian matrix in AgA, the algorithm introduces an additional computational complexity of $\mathcal{O}(n)$ to compute the Hessian-vector product. Although we strive to align the objectives of both individuals and the collective, our primary focus in this paper is on converging towards stable fixed points of the collective objective. For future research, we plan to delve deeper into the interaction of individual objectives in mixed-motive games, with an increased emphasis on understanding their dynamics in real-world mixed-motive cooperation scenarios.

### v3

@@ 1 Introduction
Recent studies in mixed-motive cooperative MARL have largely employed hand-crafted designs to promote collaboration. One popular approach is to align objectives as per existing mechanisms in cooperative games, such as reputation (Anastassacos et al., 2021), norms (Vinitsky et al., 2023), and contracts (Hughes et al., 2020). Another prevalent method leverages intrinsic motivation to align individual and collective objectives, enhancing altruistic collaboration by integrating heuristic knowledge into the incentive function. Conventionally, some works accumulate individual rewards with the group to promote altruistic conduct (Hostallero et al., 2020; Peysakhovich and Lerer, 2018; Apt and Schäfer, 2014; Roesch et al., 2024). Furthermore, some studies derive more sophisticated preference signals from the rewards of other agents (Hughes et al., 2018; McKee et al., 2020a). Additionally, several approaches aim to learn the potential influences of an agent’s actions on others (Yang et al., 2020; Jaques et al., 2019; Lu et al., 2022). Most of these algorithms rely heavily on carefully crafted designs, necessitating significant human expertise and detailed domain knowledge. On the other hand, several studies leverage Nash equilibria and related game theory concepts, such as the price of anarchy, to automatically modify rewards by learning additional weights that adjust the original objectives (Gemp et al., 2020; Kwon et al., 2023). However, finding Nash equilibria in nonconvex games presents a greater challenge compared to identifying minima in neural networks (Letcher et al., 2019).
@@ 2 Related Work
Mixed-motive cooperation refers to scenarios where the group’s objectives are sometimes aligned and at other times conflicted (Philip S. Gallo and McClintock, 1965). Recently, there has been a surge in academic interest in the Sequential Social Dilemma (SSD) (Leibo et al., 2017), which expands the concept from its roots in matrix games (Macy and Flache, 2002) to Markov games. Prosocial (Peysakhovich and Lerer, 2018) improves collective performance by blending individual rewards to redraft the agent’s overall utility. Inequity aversion further integrates the concept into Markov games by adding envy (disadvantageous inequality) and guilt (advantageous inequality) rewards to the original individual rewards (Hughes et al., 2018). To further promote cooperation, the gifting mechanism—a crucial strategy in mixed-motive cooperation (Lupu and Precup, 2020)—allows agents to influence each other’s reward functions through peer rewarding. Besides, PED-DQN (Hostallero et al., 2020) introduces an automatic reward-shaping MARL method that gradually adjusts rewards to shift agents’ actions from their perceived equilibrium towards more cooperative outcomes. Social Value Orientation (SVO) strategy introduces a unique shared rewards-based compensation approach, encouraging behavior modifications in line with interdependence theory (McKee et al., 2020b). The LIO strategy bypasses the need to modify extrinsic rewards by empowering an agent to directly influence its partner’s actions (Yang et al., 2020). Meanwhile, other studies introduce new cooperative mechanisms such as incorporating a reputation model (McKee et al., 2023), social norms (Vinitsky et al., 2023). Recently, Roesch et al. (2024) proposed a selfishness level method that incorporates social welfare into individual rewards to enhance altruistic cooperation. Additionally, several works have explored automatically modifying rewards online. D3C (Gemp et al., 2020) learns to mix rewards to improve efficiency in a Nash equilibrium. Kwon et al. (2023) extended D3C by addressing the problem of automatically modifying individual agent objectives to optimize a desired global objective. Despite these advances, many current studies lack both cost efficiency in design and theoretical analysis of alignment and convergence. To address this, our study applies a gradient perspective to align goals and further investigates the learning dynamics of our proposed method.
@@ 4.1 Differentiable Mixed-motive Game
Consider a two-player DMG with $\ell_{1}(a_{1},a_{2})=-sin(a_{1}a_{2}+a_{2}^{2})$ and $\ell_{2}(a_{1},a_{2})=-[cos(1+a_{1}-(1+a_{2})^{2})+a_{1}a_{2}^{2}]$, where $a_{i}$ represents the action of the player $i$ ($i=1,2$), and $a_{i}\in\mathbb{R}$. The rewards for the two players are the negation of their respective losses.
@@ 4.2 Altruistic Gradient Adjustment
Definition 4.2 (Altruistic Gradient Adjustment).
@@
Note that the Hessian matrix ${\bm{H}}_{c}$ is symmetric. For brevity, we denote $\nabla{\mathcal{H}}_{c}$ as $\nabla_{\bm{w}}\left(\frac{1}{2}\|{\bm{\xi}}_{c}\|^{2}\right)$, which simplifies to ${\bm{H}}_{c}^{T}{\bm{\xi}}_{c}$. In our AgA method, the formula $\xi_{c}+\lambda(\xi+{\bm{H}}^{T}_{c}\xi_{c})$ includes the Hessian matrix, but we do not compute it directly. Instead, we calculate Hessian-vector products ${\bm{H}}^{T}_{c}\xi_{c}$, which suffice for determining the adjustment gradient and the sign of $\lambda$. This approach reduces computational complexity, enhancing the effectiveness of the AgA method. The cost for computing Hessian-vector products ${\bm{H}}^{T}_{c}\xi_{c}$ is $\mathcal{O}(n)$ for $n$ weights (Pearlmutter, 1994). Refer to Section 6 for a detailed analysis.
@@ 5 Experiments
To assess the effectiveness of the proposed AgA algorithm, we perform comprehensive experiments across various environments, ranging from simple to complex, and from small to large scales. The initial environment involves a two-player public goods game (see Section 5.1), followed by sequential social dilemma environments (see Section 5.2): Cleanup and Harvest, involving 5 homogeneous players. The final test is conducted in our specially developed selfish-MMM2 environment (see Section 5.3), which is both more complex and larger in scale, featuring approximately 10 controlled agents of 3 distinct types, competing against 11 adversaries. Comparative experiments employ commonly used baseline approaches, such as simultaneous optimization with individual and collective losses (termed Simul-Ind and Siml-Co), CGA (Mescheder et al., 2017), SGA (Balduzzi et al., 2018), SVO (McKee et al., 2020a), and the selfishness level approach (termed SL) (Roesch et al., 2024).
@@ 5.1 Two-Player Public Goods Game
Table 1: The comparison of the average individual rewards (denoted as $r_{1},r_{2}$), social welfare (denoted as $SW$), and equality metric (denoted as $E$) on two- player public goods game. We show the mean of value and 95% confidence interval utilizing 50 random runs.
@@
$\mathbf{E}$ 0.923$\pm$ 0.014 0.929$\pm$ 0.012 0.924$\pm$ 0.013 0.930 $\pm$ 0.011 0.941 $\pm$ 0.014 0.940 $\pm$ 0.011 0.960 $\pm$ 0.008
Results. Table 1 presents a comparison of individual rewards and the collective outcome, derived from 50 random trials, each limited to 100 steps. The rows $r_{1}$, $r_{2}$, $SW$, and $E$ represent the individual rewards for each player, the group’s total welfare (SW), and the equality metric (E), respectively. The equality metric is based on the Gini coefficient ($G$(David, 1968), with $E$ defined as $E=1-G=1-\frac{2}{n^{2}\bar{p}}\sum_{i=1}^{n}i(p_{i}-\bar{p})$, where $\bar{p}$ is the mean of the ranked payoff vector $p$, and $n$ is the total number of players. A higher $E$ value indicates greater equality among the players.
The social optimum of the game occurs when all players allocate their entire budget, achieving the highest possible social welfare score of $3$. Among all algorithms, AgA achieves the closest social welfare score to this optimum, with a value of 2.90. Furthermore, AgA exhibits the greatest degree of fairness, with two players receiving nearly identical rewards and the highest equality value, in contrast to the baseline algorithms. Furthermore, Fig. 4 in Appendix D illustrates the action distributions of these algorithms, showing that AgA actions are more tightly clustered around the social optimum actions $(1,1)$. This further indicates that AgA approaches the social optimum more effectively.
5.2 Sequential Social Dilemma: Cleanup and Harvest
Table 2: The comparison of the average equality metric (denoted as $E$) on Harvest and Cleanup. We show the mean of equality value and standard deviation utilizing three random runs.
Envs Simul-Ind Simul-Co SVO CGA SL AgA
$\lambda=0.1$ $\lambda=1$ $\lambda=100$ $\lambda=1000$
Harvest 0.973 0.975 0.974 0.950 0.972 0.981 0.988 0.982 0.980
$\pm$ 0.005 $\pm$ 0.006 $\pm$ 0.007 $\pm$ 0.051 $\pm$ 0.005 $\pm$ 0.006 $\pm$ 0.003 $\pm$ 0.012 $\pm$ 0.006
Cleanup 0.841 0.948 0.902 0.903 0.946 0.940 0.956 0.959 0.905
$\pm$ 0.071 $\pm$ 0.013 $\pm$ 0.019 $\pm$ 0.034 $\pm$ 0.016 $\pm$ 0.017 $\pm$ 0.007 $\pm$ 0.011 $\pm$ 0.022
Baseline Comparison. As illustrated in Fig. 3(g) and Fig. 3(h), AgA demonstrates a notable improvement in social welfare over Simul-Ind, Simul-Co, CGA, SVO, and SL. Particularly, within the Cleanup scenario, AgA recorded an average social welfare of 105.15, marking approximately a 56% increase over the runner-up method SL. In the Harvest scenario, AgA shows an average improvement of 33.55 in social welfare over the next best method SVO. Table 2 compares the mean and standard deviation of the equality metric achieved by different methods in the Harvest and Cleanup environments. A value closer to 1 signifies more equal rewards among agents. As shown in the table, AgA outperforms the baseline methods, demonstrating its ability to effectively balance the interests of all team members.
@@ 5.3 Selfish-MMM2
6 Discussion
Computational Cost of AgA. The altruistic gradient adjustment is expressed as $\tilde{{\bm{\xi}}}={\bm{\xi}}_{c}+\lambda({\bm{\xi}}+{\bm{H}}_{c}^{T}{\bm{\xi}}_{c})$, where the key computational cost comes from the Hessian-vector product, ${\bm{H}}_{c}^{T}{\bm{\xi}}_{c}$. The cost for computing Hessian-vector products, ${\bm{H}}_{c}^{T}\xi_{c}$, is $\mathcal{O}(n)$ for $n$ weights (Pearlmutter, 1994). This introduces added complexity compared to standard methods in mixed-motive MARL, making AgA’s running time generally more than twice as long as gradient-based methods. Table 3 provides a comparison of the average running time between AgA and baseline methods in a two-player public goods game, including total duration, timesteps, time per step, and the time ratio relative to Simul-Ind. Simul-Ind, Simul-Co, and SL are standard gradient-based methods, while CGA and SGA modify gradients. Our results show that AgA takes 2-3 times longer per step but remains the most efficient, requiring only 1389 steps over 50 runs, compared to around 4000 steps for the baselines.
Table 3: Comparison of the average running time between baseline methods and AgA in the two-player public goods game, including total duration, timesteps, time per step, and the time per step ratio relative to Simul-Ind.
Metrics Simul-Ind Simul-Co SL CGA SGA AgA
Duration (ms) 1165.79 910.15 1149.97 3041.84 3007.77 1034.69
Steps 4272 3252 3887 4478 4179 1389
Step Time (ms) 0.27 0.28 0.30 0.68 0.72 0.74
Ratio 1.00 1.04 1.11 2.52 2.67 2.74
Limitations. Despite conducting a series of experiments to verify the proposed AgA algorithm, including using a new large-scale mixed-motive cooperation testbed, these environments remain somewhat removed from real-world applications. Although we strive to align the objectives of both individuals and the collective, our primary focus in this paper is on converging towards stable fixed points of the collective objective. For future research, we plan to delve deeper into the interaction of individual objectives in mixed-motive games, with an increased emphasis on understanding their dynamics in real-world mixed-motive cooperation scenarios.
@@ 7 Conclusion
Acknowledgement
This work is partially supported by National Key R&D Program of China (2022ZD0114804) and National Natural Science Foundation of China (62106141). Jianhong Wang is fully supported by the Engineering and Physical Sciences Research Council [Grant Ref: EP/Y028732/1]. The authors also thank Shuqing Shi for his kind assistance and advice.